OpenStudy (anonymous):

Help! Please Which set of integers is a Pythagorean triple? A.6, 8, 12 B.3, 5, 6 C.15, 9, 17 D.5, 12, 13

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

you can check each by stating 6^2+8^2=12^8 true or false? 3^2+5^2=6^2 true or false? 15^2+9^2=17^2 true or false? 5^2+12^2=13^2 true or false? the one that is true is going to be the answer.

OpenStudy (anonymous):

I'm confused ( i have never been so good with these )

OpenStudy (solomonzelman):

you don't have to be so good at those, let me explain what I did, Pythagorean triple means

OpenStudy (solomonzelman):

B I* I * side c I * side a I * I * I______________________*_ C A side b the formula of how all sides are proportional if you will, a^2+b^2=c^2

OpenStudy (solomonzelman):

like if side a was 3 and side b was 4, to find c, you would a^2+b^2=c^2 3^2+4^2=c62 9+16=c^2 25=c^2 c=5 or -5 a distance of a side can't be negative, so it would be just (positive) 5.

OpenStudy (anonymous):

ok thanks!

OpenStudy (solomonzelman):

So you tell me, to verify that you are correct, what choice is your final answer?